User’s Guide to the Yearly Calendar package^{∗}

Frank G. Bennett, Jr. January 24, 1998

Abstract

This package makes a yearly calendar on one side of A4.

Please see the file calguide.tex for documentation on this calendar style.

∗This file is version number 2.1.1.6. It was last revised on 1997/11/06 07:33:21.